### 15 Mar 2024 v2

# Benchmarking News Recommendation in the Era of Green AI

Abstract.
Over recent years, news recommender systems have gained significant attention in both academia and industry, emphasizing the need for a standardized benchmark to evaluate and compare the performance of these systems. Concurrently, Green AI advocates for reducing the energy consumption and environmental impact of machine learning. To address these concerns, we introduce the first Green AI benchmarking framework for news recommendation, known as GreenRec [^1] GreenRec has been published as part of the Legommenders benchmark, which is a modular framework for recommender systems. The source code and data for GreenRec can be found at https://github.com/Jyonn/Legommenders., and propose a metric for assessing the tradeoff between recommendation accuracy and efficiency. Our benchmark encompasses 30 base models and their variants, covering traditional end-to-end training paradigms as well as our proposed efficient only-encode-once (OLEO) paradigm. Through experiments consuming 2000 GPU hours, we observe that the OLEO paradigm achieves competitive accuracy compared to state-of-the-art end-to-end paradigms and delivers up to a 2992% improvement in sustainability metrics.
journalyear: 2024copyright: acmlicensedconference: Companion Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singaporebooktitle: Companion Proceedings of the ACM Web Conference 2024 (WWW ’24 Companion), May 13–17, 2024, Singapore, Singaporedoi: 10.1145/3589335.3651472isbn: 979-8-4007-0172-6/24/05ccs: Information systems Recommender systemsccs: Information systems Data mining
1. Introduction
News recommender systems have become a crucial tool in the digital age, helping users discover relevant news articles and stay informed amidst the ever-increasing number of sources available. To integrate news content knowledge into recommender systems, content encoders are employed to capture semantic news representations. Various approaches have been explored in these efforts, including simple content encoders represented by CNN (Kim, 2014) and Attention-based (Vaswani et al., 2017) models (An et al., 2019; Wu et al., 2019a, c; Li et al., 2022), as well as emerging content encoders based on pretrained language models (Zhang et al., 2021; Liu et al., 2022).
Despite the success achieved by these studies, there is still a notable absence of standardized benchmarks and uniform evaluation protocols for news recommendation. Consequently, even when common datasets like MIND (Wu et al., 2020) are employed for evaluation, existing studies often use their own data partitions and apply unique preprocessing steps. The absence of uniformity in data preprocessing methods leads to non-reproducible and often conflicting experimental outcomes across different studies, and the use of non-standard data preprocessing techniques makes it challenging to compare and evaluate results between papers.
Meanwhile, the environmental impact has also begun to draw attention from the machine learning community, as training large-scale networks leads to a huge increase of carbon emission (Spillo et al., 2023). Therefore, a series of studies (Wilson et al., 2020; Przybyła and Shardlow, 2022) explore ways to reduce energy consumption in line with Green AI principles (Schwartz et al., 2020).
Figure 1. Two training paradigms for news recommendation.
In light of this, we establish the first Green AI Benchmark for News Recommendation, called GreenRec, and introduce a sustainability metric, namely unit conversion rate of carbon emission, to strike a balance between recommendation quality and efficiency. Specifically, we selected six base models and considered five different variants for each model. In total, we evaluated the recommendation accuracy and environmental impact across a total of 30 baselines. This evaluation was carried out under two distinct training paradigms: the conventional end-to-end training paradigm, where the content encoder and other recommender components are jointly trained, and our proposed only-encode-once (OLEO) training paradigm, where the content encoder and other components are trained in a decoupled manner. Our evaluation is conducted through hundreds of experiments within a standardized framework using the most widely used news recommendation datasets, MIND-small and MIND-large (Wu et al., 2020). The experiments reveal that the conventional end-to-end paradigm encounters efficiency challenges, while our OLEO-based paradigm is much more eco-friendly. Moreover, the OLEO-based model variants maintain competitive performance with the state-of-the-art end-to-end PLM-NR variants (Wu et al., 2021), and achieve up to 2992% improvement in our sustainability metric.
2. News Recommendation
2.1. Basics of News Recommender Systems
News recommender systems are devised to predict the click probability of a user over a piece of given news article. Typically, a general news recommender system consists of three main components: content encoder, history encoder, and interaction module. As depicted in Figure 1 (a), the recommender accepts a news-user pair as input and yields the click probability. The content encoder is responsible for capturing the semantic meaning and contextual information of the news articles, while the history encoder processes the news history vectors, capturing the preferences, interests, and browsing patterns of users. Finally, the interaction module facilitates the interaction between the candidate news vector and user vector to calculate the click probability for the given news-user pair.
2.2. Training Paradigms and Complexity
Traditional models (Wu et al., 2019b, a, c, 2021) typically follow such end-to-end training paradigm, as displayed in Figure 1 (a), where the three components are are trained together, guided by click-through data.
However, this traditional approach has shown limitations, particularly in terms of efficiency. A significant issue is the repetitive work done by the content encoder due to the repeated encoding of the same news articles, a consequence of their frequent interaction with various users. For instance, if there are $N$ news articles, $M$ users, and $K$ instances of user-news interactions, with an average historical sequence of $L$ interactions per user, the content encoder ends up processing $M\times L$ articles. The overall time required for training can be roughly calculated as $O(M\times L\times t_{c}+K\times(t_{u}+t_{i}))$, highlighting a particularly heavy load when $M\times L$ significantly exceeds either $N$ or $K$, especially with advanced content encoders (Zhang et al., 2021; Wu et al., 2021) that have a much higher time complexity compared to other modules. For example, in the MIND dataset scenario, a single news article is encoded on average about 1818 times in one epoch of training, indicating a substantial amount of redundancy.
To address these inefficiencies, we draw inspiration from a new training approach known as the only-encode-once paradigm, as depicted in Figure 1 (b). This approach involves initially pretraining the content encoder using the news content in a self-supervised manner. After this phase, the content vectors are extracted and stored, allowing the history encoder and interaction module to be trained separately with much quicker access to the news representations via a lookup table. The pretraining phase for the content encoder has a time complexity of $O(N\times t_{c})$, considered a fixed cost across various models. Meanwhile, the training of the recommendation system itself becomes significantly more efficient, with a time complexity of $O(K\times(t_{u}+t_{i}))$, marking a drastic reduction from the traditional method.
Usage TorchRec DeepCTR DeepRec RecBole FuxiCTR BARS Ours
Data Preprocessing
Content-aware CTR $\times$ $\times$ $\times$
NewsRec Methods $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Decoupled Training $\times$ $\times$ $\times$
End-to-end Training $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Green AI Evaluation $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Table 1. Comparison to exisiting recommendation benchmarks ($\times$ means totally — partially — not met, respectively). “Partially met” represents incomplete availability.
2.3. Adapting CTR Models to both Paradigms
While the prevailing news recommendation models adhere to a matching-based approach with simple dot product as the interaction module, preserving all positive samples and conducting negative sampling (including real or fake negatives) and trained as a classification task, there are still some studies (Zhang et al., 2021; Liu et al., 2022) that have validated the effectiveness of ranking-based CTR methods in news recommendation systems. In these approaches, all samples are retained throughout training, and the click probability is directly produced as a regression task. These CTR techniques incorporate intricate feature interaction networks, which can be utilized in the interaction module. When integrated into the content encoder, either through the end-to-end training paradigm or by initializing with pretrained news representations in the OLEO training paradigm, they gain the content-aware abilities.
2.4. Proposed GreenRec Benchmark
Here, we summarize 6 representative methods and 5 variants used in our GreenRec benchmark:
Base models. We select three representative news recommenders (i.e., NAML (Wu et al., 2019a), LSTUR (An et al., 2019), and NRMS (Wu et al., 2019c)) and three widely used CTR methods (i.e., BST (Chen et al., 2019), DCN (Wang et al., 2017), and DIN (Zhou et al., 2018))
ID-based variants. [^2] We use green, yellow, and red text to represent OLEO-based or content-free (ID-based) training methods, end-to-end training variants with simple content encoders and end-to-end training variants with pretrained language models. We remove the content encoder from the news recommenders and replace it with a randomly initialized embedding lookup table. CTR methods are inherently ID-based.
Text-based variants. We remove the news embedding lookup table of the CTR methods and replace it with a content sequence pooling layer. News recommenders are inherently text-based.
PLM-NR variants. As proposed by PLM-NR (Wu et al., 2021), we replace the content encoders of the text-based variants with pretrained language models such as BERT (Devlin et al., 2018).
PREC variants. As proposed by PREC (Liu et al., 2022), we employ the OLEO training paradigm which first pretrains the content encoder and then replace the content encoders of the text-based variants with the informative news embedding lookup table.
BERT-OLEO variants. Unlike the PREC variants which requires the pretraining phase, here we directly use the pretrained BERT model to extract and cache the news representations, and construct the informative lookup table.
3. Implementation and Evaluation
In this section, we first compare our GreenRec with other benchmarking studies. Then, we outline our evaluation procedures, including quality-based assessments and our proposed sustainability metric. Finally, we report and discuss our benchmarking results.
3.1. Comparison with Previous Benchmarks
In recent years, the issues of non-reproducibility and unfair comparisons have garnered increasing attention within the recommendation community. A series of studies (Ivchenko et al., 2022; Shen, 2017; Zhang et al., 2019; Zhao et al., 2022; Zhu et al., 2021, 2022) have conducted comprehensive experiments and analyses, advocating for a unified benchmark construction. As shown in Table 1, we compare our news recommendation benchmark GreenRec with these existing efforts. It is evident that these studies primarily target in the general recommendation domain, particularly emphasizing aspects such as CTR prediction and collaborative filtering. Nevertheless, a standardized benchmarking pipeline tailored to the news recommendation domain and the environmental evaluation remains conspicuously absent. Hence, our GreenRec is the first attempt on benchmarking the news recommenders in the era of Green AI.
3.2. Dataset and Data Splitting
##\## News ##\## Users ##\## Interactions ##\## Samples Density
MIND-small 65,238 94,057 347,727 8,381,093 0.1366%
MIND-large 104,151 750,434 3,958,501 95,447,571 0.1221%
Table 2. Dataset statistics. Density is defined as the ratio of # interactions to # all possible interactions.
Dataset MIND-small MIND-large
Method Matching CTR Matching CTR
NAML LSTUR NRMS BST DCN DIN NAML LSTUR NRMS BST DCN DIN
ID-based AUC 50.13 51.04 54.84 50.09 53.92 55.95 52.98 54.98 57.59 52.10 57.41 57.36
MRR 23.01 22.90 26.53 22.13 25.18 25.88 24.52 25.99 27.41 24.81 26.76 26.70
N@5 22.35 22.31 26.34 21.59 24.43 25.95 24.12 25.64 27.05 24.63 26.90 26.84
[…]
ApC 0.68 5.20 17.29 0.24 6.53 7.08 1.01 1.41 1.61 0.38 0.80 0.57
Text-based (End-to-end) AUC 60.14 61.27 62.21 60.51 62.63 62.90 63.03 63.89 64.12 63.28 63.88 64.02
MRR 28.93 29.64 30.19 28.59 29.73 30.06 30.40 31.24 31.77 30.73 31.65 31.98
N@5 29.33 30.28 31.10 29.09 30.52 30.65 31.82 32.15 32.64 31.95 32.40 33.00
[…]
ApC 24.14 19.43 19.69 19.83 20.05 14.33 2.01 1.56 1.40 1.09 1.43 1.01
PLM-NR (End-to-end) AUC 62.06 63.64 62.53 64.40 63.32 63.26 65.19 65.73 65.57 66.03 65.42 65.31
MRR 31.66 31.74 30.74 32.21 32.00 31.83 32.74 33.18 32.94 33.40 32.85 32.68
N@5 32.25 32.72 31.31 33.34 32.58 32.40 33.77 34.26 34.13 34.70 33.99 33.70
[…]
ApC 6.78 6.75 4.97 2.85 0.76 0.72 0.60 0.52 0.39 0.20 0.06 0.05
BERT (OLEO) AUC 60.62 61.09 60.94 60.81 62.65 62.40 63.02 63.62 63.40 62.94 64.29 63.75
MRR 29.31 29.26 29.31 29.04 30.92 30.75 31.23 31.59 31.38 30.56 32.60 31.58
N@5 29.71 29.60 29.65 29.38 31.37 32.44 31.79 32.30 32.16 31.83 33.63 32.43
[…]
ApC 48.27 48.22 33.15 28.45 20.40 14.41 3.69 3.37 2.65 2.02 1.49 1.44
PREC (OLEO) AUC 62.95 62.16 62.95 62.43 64.57 63.12 64.78 64.88 64.34 65.33 65.44 64.53
MRR 31.26 31.00 31.18 30.42 32.60 31.28 32.64 32.94 32.93 33.29 33.04 32.72
N@5 32.01 31.79 32.10 30.94 33.48 32.01 33.66 34.00 33.95 34.35 34.03 33.58
[…]
ApC 58.86 52.87 39.24 32.71 23.50 15.25 4.19 3.68 2.84 2.40 1.61 1.52
ApC Imp. (%) 768% 683% 690% 1048% 2992% 2018% 598% 608% 628% 1100% 2583% 2940%
Table 3. Comparison of variants of different recommenders. “ApC Imp” represents the environmental sustainability growth rate of PREC variants compared to the SOTA PLM-NR variants. We bold the best results and underline the second-best results.
In this work, we mainly use two versions of the most prevalent news recommendation dataset: MIND-small and MIND-large (Wu et al., 2020). Table 2 summarizes the data statistics information.
Due to the fact that the MIND-small dataset only provides training and validation sets, many studies have failed to specify their exact partitioning during training, and in some cases, they have directly reported results on the validation set, which leads to an unfair comparison among different models. Furthermore, the MIND-large dataset only offers online testing, which is not conducive to offline environment assessment. To address these issues, we have restructured both datasets by splitting the validation sets of both datasets in a 1:1 ratio to create new validation and testing sets. The scripts can be found in our released repository.
3.3. Evaluation Metrics
[…]
(1) $CO_{2}E=p\times t\times c,$
where $p$ is the power consumption differ by hardware types, $t$ is the hardware running time, and $c$ is carbon efficiency differ by power provider. ApC measures the unit conversion rate of carbon emissions on AUC, defined as:
(2) $ApC=\frac{AUC-50}{CO_{2}E}\times 100.$
Training Details. The Adam optimizer is employed for optimization. We consider various learning rates from the set $\{1e-5,2e-5,5e-5,1e-4,2e-4,5e-4\}$ and batch sizes from the set
$\{64,128,256,500,1000,5000\}$. The embedding dimensions for all base models and their variants are fixed to 64. For the OLEO training paradigm which uses cached news vectors for initialization, we keep the pretrained embeddings fixed and use learnable transformation matrices for dimension projection. Experimental results are averaged over five runs, and all methods were trained using Nvidia GeForce RTX 3090 with 24GB memory, which has a power consumption of 350W. The carbon efficiency in our experimental region (anonymous for review) is 722g CO2-eq/kWh.
3.4. Findings
The comparative results of various models are summarized in Table 3, from which we derive the following observations.
ID- vs. Modality-based Recommender. Among the five variants, the ID-based variants exhibit the worst performance, showing the significance of news content comprehension in news recommenders. The findings in recent study (Yuan et al., 2023) are in line with ours.
Enhancement of PLMs in the end-to-end training paradigm. The PLM-NR variants consistently outperforms the ID-based and Text-based ones, benefiting from the augmented understanding of news content offered by pretrained knowledge. However, they grapple with an efficiency issue, resulting in significantly higher carbon emissions compared to other variants.
Enhancement of PLMs in the OLEO training paradigm. The BERT variants maintain an environmentally sustainable level of efficiency. Nonetheless, there is a deficiency in the ability of general language models to fully grasp news content, resulting in a notable decline in accuracy compared to the PLM-NR variants. However, PREC variants achieve comparable performance to PLM-NR ones while significantly reducing carbon emissions, which represents the prime case of the tradeoff between accuracy and efficiency.
4. Conclusion
We have established the first standardized Green AI benchmark for news recommendation. Through extensive experiments over 30 base models and their variants, we conclude that the OLEO training paradigm successfully strikes a balance between accuracy and efficiency. We encourage other researchers to explore the OLEO paradigm and make contributions to environmental conservation.
Acknowledgement
Qijiong Liu would like to thank Dr. Guanliang Chen and Mr. Haoxiang Shi for their insightful input during the brainstorming process of our paper. Additionally, the authors would like to thank the anonymous reviewers for their valuable feedback.